### v1

@@
11institutetext: Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Straße 1, 85748 Garching bei München, Germany
11email: korol@mpa-garching.mpg.de
22institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, UK, NE1 7RU
22email: andrei.igoshev@newcastle.ac.uk
Probing the millisecond pulsar origin of the GeV excess
@@
Valeriya Korol, These authors contributed equally to this work.11 0000-0002-6725-5935    Andrei Igoshev, These authors contributed equally to this work.22 0000-0003-2145-1022
(Received September $X$, 2025; accepted September $Y$, 2025)
The GeV $\gamma$-ray excess observed towards the Galactic Centre remains unexplained. While dark matter annihilation has long been considered a leading interpretation, an alternative scenario involving a large population of millisecond pulsars has not been ruled out. Testing this hypothesis with electromagnetic observations is difficult, as pulsar searches in the bulge are strongly affected by scattering, high sky temperature, and source confusion. We investigate whether gravitational-wave observations with the Laser Interferometer Space Antenna (LISA) could provide an independent probe of the millisecond pulsar binary population in the Galactic bulge.
@@
LISA will measure binary frequencies with high precision, but chirp masses can only be determined for the most massive or highest-frequency systems. Distinguishing millisecond pulsar binaries from the far more numerous double white dwarfs will be challenging, though LISA detections could provide valuable targets for follow-up with the Square Kilometre Array, enabling a critical test of the millisecond pulsar origin of the $\gamma$-ray excess.
@@
gravitational waves – binaries: close – pulsars: general – white dwarfs – Galaxy: bulge – gamma rays: diffuse background
@@ 1 Introduction
More than a decade ago, an excess of $\gamma$-ray emission from the direction of the Galactic Centre was discovered in data from the Fermi Gamma-ray Space Telescope (Goodenough & Hooper, 2009; Hooper & Goodenough, 2011; Hooper & Linden, 2011). Multiple studies independently discovered an excess of $\gamma$-ray emission in comparison to state-of-the-art background modelling peaking around 1 – 3 GeV. This emission extends to at least $10^{\circ}$ from the centre and with approximate spherical symmetry. The emitting region roughly overlaps with the Galactic bulge and has a spatial extent of $\approx 3$ kpc. Significant efforts have been made to test the robustness of the $\gamma$-ray background models (e.g., Boyarsky et al. 2011; Gordon & Mac´ıas 2013; Calore et al. 2015), and the GeV excess has withstood these tests. The leading explanation proposed for this excess, suggested by Goodenough & Hooper (2009), is that the excess originates from dark matter annihilation.
@@
A few years prior to the discovery of the GeV excess, Baltz et al. (2007) realised that high-energy emission from radio pulsars can mimic the spectral properties of dark matter annihilation. Because pulsars produce high-energy emission with roughly similar spectra peaking around 2 GeV, Hooper & Goodenough (2011) put forward the idea that a large population of unresolved (millisecond) pulsars could be the astrophysical source of the $\gamma$-ray excess. This hypothesis has received significant support over the years (Abazajian, 2011; Mirabal, 2013; Bartels et al., 2016).
@@
Millisecond radio pulsars (MSPs; see Lorimer 2008 for a review), also known as recycled pulsars, are neutron stars (NSs) characterised by short spin periods ($P_{\rm s}<30$ ms) and small spin period derivatives ($\dot{P}_{\rm s}<10^{-19}$ s s-1). They are thought to acquire these properties through stable mass transfer in low-mass X-ray binaries (LMXBs), often accreting from a white dwarf (WD) companion. During this accretion phase, the NS gains angular momentum and is spun up to millisecond periods—a process known as the recycling mechanism (Alpar et al., 1982; Bhattacharya & van den Heuvel, 1991). The accreted material also buries the NS’s magnetic field, reducing it to typical values of $^{8}1010$$^{10}$ G. Once accretion ceases, the NS begins to lose rotational energy through magnetic dipole radiation, becoming a $\gamma$-ray source. The $\gamma$-ray luminosity is proportional to the spin-down luminosity ($\dot{E}\propto\dot{P}_{\rm s}/P_{\rm s}^{3}$), meaning that short spin periods result in strong $\gamma$-ray emission. Despite their relatively weak dipolar fields, MSPs are among the brightest $\gamma$-ray sources in the sky. As the pulsar ages, its period gradually increases over gigayear timescales, leading to a slow decline in $\gamma$-ray luminosity.
@@
Detecting MSPs in the Galactic Centre through radio observations has proven challenging. The region is heavily affected by interstellar scattering and dispersion, which smear out the rapid pulsations of MSPs and reduce their detectability in standard radio surveys (e.g. Cordes & Lazio, 1997). In addition, the high sky temperature and source confusion in the dense stellar environment further complicate searches. As a result, the true size of the MSP population in this region remains unknown. To date, only seven radio pulsars have been discovered in its vicinity (e.g. Johnston et al., 2006; Deneva et al., 2009; Bates et al., 2011; Eatough et al., 2013; Wongphechauxsorn et al., 2024). Zhao et al. (2020) identified 110 compact radio sources, and Zhao et al. (2022) discovered 64 hypercompact sources in the same area, some of which may be radio pulsars. Multi-wavelength analyses have provided indirect constraints on the pulsar population. For instance, Wharton et al. (2012) estimated that the inner parsec (i.e. a region $\approx$300 times smaller than the $\gamma$-ray excess area) around the Galactic Centre could host up to $10^{3}$ active pulsars. Over the years, several studies have attempted to estimate the size of the MSP population that could be responsible for $\gamma$-ray excess (Hooper & Linden, 2016; Ploeg et al., 2017; Dinsmore & Slatyer, 2022). A recent study by Holst & Hooper (2025) derived a $\gamma$-ray luminosity function using field MSPs and concluded that, if $\gamma$-ray luminosity remains constant over time, the Galactic bulge could host as many as $3.6\times 10^{4}$ MSPs. If instead the luminosity decays with time, the estimated population size rises to $2\times 10^{5}$ MSPs.
@@
NSs/MSPs are also promising sources of continuous gravitational wave emission. If a rapidly spinning NS has even a small amount of ellipticity—either imprinted at birth, caused by starquakes, accretion from a companion, or sustained by internal magnetic fields—it will be characterised by a non-zero mass quadrupole. This leads to the emission of gravitational waves at a frequency twice the star’s rotational frequency, typically in the tens to thousands of hertz range, which falls within the sensitivity band of ground-based detectors (Friedman & Schutz, 1978). However, these continuous signals are intrinsically weak, and at the distance of the Galactic Center, they remain well below the detection threshold of current detectors (e.g., Abbott et al., 2022; Riles, 2023; Steltner et al., 2023). Next-generation detectors, such as the Einstein Telescope and Cosmic Explorer, are expected to achieve the sensitivity required to probe these sources in the Galactic Center (e.g., Pagliaro et al., 2023; Bartel & Profumo, 2024).
@@
If the orbital period is shorter than about two hours, the signal falls in the millihertz regime, placing it within reach of space-based detectors such as the Laser Interferometer Space Antenna (LISA Colpi et al., 2024), even at distances comparable to the Galactic Centre and beyond (Seto, 2019; Lau et al., 2020a; Korol & Safarzadeh, 2021; Wagg et al., 2022). In this article, we ask whether LISA can probe the MSP population in the Galactic bulge and thus provide an independent test of their proposed contribution to the $\gamma$-ray excess. We argue that, because gravitational-wave detection is subject to entirely different selection effects than electromagnetic searches for MSPs in the Galactic Centre, LISA offers a promising new way to test the MSP origin of the $\gamma$-ray excess. To illustrate this potential, we present the first forecasts for MSP binaries in the Galactic bulge detectable by LISA.
@@
This article is structured as follows. In Section 2, we briefly review the literature on the formation of MSP binaries and summarise their gravitational-wave detectability with LISA. In Section 3, we describe the assumptions and methodology behind our population synthesis of the bulge. In Section 4, we present our forecasts for different bulge formation scenarios and discuss the likelihood of detecting and characterising the MSP population with LISA. Finally, in Section 5, we interpret our findings in a broader context, and summarise our conclusions in Section 6.
@@ 2 Galactic millisecond pulsar binary population
The Australia Telescope National Facility (ATNF) pulsar catalogue v 2.6.1111https://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005) contains information on 624 pulsars with spin periods shorter than $P\leq 30$ ms, most of which are classified as MSPs. A significant fraction of these pulsars are located in globular clusters. For example, the most recent compilation by Paulo Freire [^2] https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html summarises properties of 345 pulsars in 45 clusters, indicating that at least 55% of all known MSPs reside in globular clusters. Whether formed dynamically in globular clusters or through isolated binary evolution in the field, MSPs can end up in a variety of binary configurations. Below, we introduce those configurations that are mentioned in this study, drawing primarily from the classification and discussion presented in Tauris & van den Heuvel (2023).
@@
Many MSPs are thought to descend from LMXBs, in which a NS accretes material from a companion star and is spun up to millisecond periods. A large fraction of known MSPs are found in binaries with helium-core WD (He WD) companions. Other systems host more massive carbon–oxygen-core WDs (CO WDs) or oxygen–neon–magnesium-core WDs (ONeMg WDs). Some MSPs reside in tight ultra-compact X-ray binaries (UCXBs) with orbital periods as low as a few minutes, while others are found in eclipsing systems known as spiders. These are typically categorised as black widows, with extremely low-mass companions (<0.1<0.1 M), or redbacks, with more substantial companions (0.10.10.40.4 M). Most recently, Yang et al. (2025) discovered an MSP with a massive companion (up to $1.6$ M), most likely a stripped helium star (He star).
@@
Figure 1: Top panel: Orbital periods and companion masses for MSPs compiled from the ATNF Pulsar Catalogue v2.6.0. The companion mass shown is the median mass, computed assuming an orbital inclination of $i=60^{\circ}$. All data points (blue cross symbols) correspond to radio-detected MSPs. Among them, $\gamma$-ray sources are highlighted with green downward-pointing triangles, and those located in globular clusters (GC) are marked with orange upward-pointing triangles. Grey error bars indicate lower and upper limits on WD companion masses. For reference, the theoretical orbital period–companion mass relation from Tauris & Savonije (1999) for Population I stars is shown as a solid red line. The minimum gravitational-wave frequency detectable by LISA is shown as a dashed black line.
@@
Bottom left panel: Distribution of companion masses for the MSP sample. The solid red lines show a power-law distribution with a slope of 4-4, plotted for comparison. Bottom right panel: Distribution of binary orbital periods for the MSP sample. Overplotted are the estimated pulsar detection efficiencies in radio surveys, based on the modelling by Bagchi et al. (2013). Two search strategies are shown: $\gamma_{1}$ for a standard (non-accelerated) search, and $\gamma_{2}$ for an acceleration search. These detection efficiency curves are computed assuming the following orbital parameters: eccentricity $e=0$, NS mass $M_{\mathrm{MSP}}=1.4$ M, integration time $T\mathrm{obs}=1000$ s, WD companion mass derived from the Tauris & Savonije (1999) relation, spin period of 3 ms, orbital inclination $i=60^{\circ}$, and longitude of periastron $\Omega=30^{\circ}$.
@@ 2.1 MSPs formation and orbital parameters
Tauris & Savonije (1999) identified a relation between the mass of a He WD and the orbital period at the end of stable mass transfer. This relation arises because stable mass transfer during the LMXB stage uniquely determines the orbital separation based on the donor star’s density. The relation is given by:
@@
To better understand the observed MSP population, we select all millisecond radio pulsars from the ATNF pulsar catalogue. We select sources using the commonly adopted definition with $P_{\rm s}<30$ ms and $\dot{P}_{\rm s}<10^{-19}$ s/s. We also retrieve information about associations with globular clusters and the presence of $\gamma$-ray emission. The resulting distribution of orbital periods and companion masses is shown in the top panel of Figure 1. The MSPs fall broadly into two groups: (1) binaries with He WD companions that follow the Tauris & Savonije (1999) relation (red solid line), and (2) binaries with low-mass companions—including systems that are the products of UCXB evolution, redbacks, and black widows. The ATNF catalogue also includes some MSPs likely formed through a common-envelope channel with CO WD companions, as well as rare systems in which MSPs are paired with NSs (i.e. those with companion masses >0.5>0.5 M in Figure 1). Importantly, these latter systems are only partially recycled, meaning they have longer spin periods and higher magnetic fields. As a result, their spin-down luminosities are typically too low to produce detectable $\gamma$-ray emission, and these pulsars are therefore not seen as $\gamma$-ray sources. Nevertheless, some of them may be prominent gravitational-wave sources because of their larger component masses.
In the bottom panels of Figure 1, we show the distributions of companion masses (left) and orbital periods (right) for the MSPs in our ATNF catalogue selection. The companion mass distribution can be approximated by a power law with an exponent of 4-4, which matches the observed trend down to $\sim 0.2$ M, particularly for systems in globular clusters (orange distribution). In the right panel, we compare the orbital period distribution with the detection efficiency curves from Bagchi et al. (2013) for standard ($\gamma_{1}$) and acceleration ($\gamma_{2}$) radio searches. The number of known MSPs drops sharply at short orbital periods, closely following the shape of the $\gamma_{1}$ efficiency curve. This suggests that observational biases significantly affect the detection of binaries with orbital periods shorter than one day, and implies that the intrinsic population may be more uniformly distributed in logarithmic orbital period space than the observed sample indicates.

### v2

@@
11institutetext: Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Straße 1, 85748 Garching bei München, Germany 22institutetext: SRON, Netherlands Institute for Space Research, Niels Bohrweg 4, 2333 CA Leiden, The Netherlands
22email: v.korol@sron.nl
33institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, UK, NE1 7RU
33email: andrei.igoshev@newcastle.ac.uk
Probing the millisecond pulsar origin of the $\gamma$-ray excess
@@
(Received: 16 September 2025 — Accepted: 8 November 2025)
The gigaelectronvolt $\gamma$-ray excess observed towards the Galactic centre remains unexplained. While dark matter annihilation has long been considered a leading explanation, an alternative scenario involving a large population of millisecond pulsars remains viable. Testing this hypothesis with electromagnetic observations is difficult, as pulsar searches in the bulge are strongly affected by interstellar scattering, high sky temperature, and source confusion. We investigate whether gravitational-wave observations with the Laser Interferometer Space Antenna (LISA) could provide an independent probe of the millisecond pulsar binary population in the Galactic bulge in the future.
We constructed synthetic populations of detached millisecond pulsar–white dwarf binaries under two illustrative formation scenarios: an accreted scenario, in which systems are deposited by disrupted globular clusters, and an in situ scenario, in which binaries form through isolated binary evolution. In both cases, only $10^{-5}$$10^{-4}$ of the underlying bulge population is detectable by LISA. Still, even a few detections would imply tens to hundreds of thousands of unseen systems. Accreted binaries are expected to have lower chirp masses ($\sim$0.4 M), while in situ binaries produce more massive companions ($\sim$0.9 M), though part of this contrast reflects our modelling assumptions. LISA will measure binary frequencies with high precision, but chirp masses can only be determined for the most massive or highest-frequency systems. Thus, identifying millisecond-pulsar binaries among the far more numerous double white dwarfs will be challenging, as their gravitational-wave signals alone are indistinguishable. However, coordinated follow-up with the Square Kilometre Array of LISA-selected targets could directly test the millisecond-pulsar explanation of the $\gamma$-ray excess.
@@
gravitational waves – binaries: close – pulsars: general – Galaxy: bulge – gamma rays: diffuse background
@@ 1 Introduction
More than a decade ago, an excess of $\gamma$-ray emission from the direction of the Galactic centre was discovered in data from the Fermi Gamma-ray Space Telescope (Goodenough and Hooper, 2009; Hooper and Goodenough, 2011; Hooper and Linden, 2011). Multiple studies independently confirmed an excess of $\gamma$-ray emission, relative to state-of-the-art background modelling, peaking around 1–3 GeV. This emission extends to at least $10^{\circ}$ from the centre and exhibits approximately spherical symmetry. The emitting region roughly overlaps with the Galactic bulge and has a spatial extent of $\approx 3$ kpc. Significant efforts have been made to test the robustness of the $\gamma$-ray background models (e.g. Boyarsky et al. 2011; Gordon and Macías 2013; Calore et al. 2015; Zhou et al. 2015) and the gigaelectronvolt excess has withstood these tests. The leading explanation for this excess, first proposed by Goodenough and Hooper (2009), is that it originates from dark matter annihilation.
@@
A few years prior to the discovery of the gigaelectronvolt excess, Baltz et al. (2007) realised that high-energy emission from radio pulsars can mimic the spectral properties of dark matter annihilation. Because pulsars produce high-energy emission with roughly similar spectra peaking around 2 GeV, Hooper and Goodenough (2011) put forward the idea that a large population of unresolved (millisecond) pulsars could be the astrophysical source of the $\gamma$-ray excess. This hypothesis has received significant support over the years (Abazajian, 2011; Mirabal, 2013; Bartels et al., 2016).
@@
Millisecond radio pulsars (MSPs; see Lorimer 2008 for a review), also known as recycled pulsars, are neutron stars (NSs) characterised by short spin periods ($P_{\rm s}<30$ ms) and small spin period derivatives ($\dot{P}_{\rm s}<10^{-19}$ s s-1). They are thought to acquire these properties through stable mass transfer in low-mass X-ray binaries (LMXBs). During this accretion phase, the NS gains angular momentum and is spun up to millisecond periods—a process known as the recycling mechanism (Alpar et al., 1982; Bhattacharya and van den Heuvel, 1991). The accreted material also buries the NS’s magnetic field, reducing it to values typically of $^{8}1010$$^{10}$ G. Once accretion ceases, the NS begins to lose rotational energy through magnetic dipole radiation, becoming a $\gamma$-ray source. The $\gamma$-ray luminosity is proportional to the spin-down luminosity ($\dot{E}\propto\dot{P}_{\rm s}/P_{\rm s}^{3}$), meaning that short spin periods result in strong $\gamma$-ray emission. Despite their relatively weak dipolar fields, MSPs are among the brightest $\gamma$-ray sources in the sky. As the MSP ages, its spin period gradually increases over gigayear timescales, leading to a slow decline in $\gamma$-ray luminosity.
@@
Detecting MSPs in the Galactic centre through radio observations has proven to be challenging. The region is heavily affected by interstellar scattering and dispersion, which smear out the rapid pulsations of MSPs and reduce their detectability in standard radio surveys (e.g. Cordes and Lazio, 1997). In addition, the high sky temperature and source confusion in the dense stellar environment further complicate searches. As a result, the true size of the MSP population in this region remains unknown. To date, only seven radio pulsars have been discovered in its vicinity (e.g. Johnston et al., 2006; Deneva et al., 2009; Bates et al., 2011; Eatough et al., 2013; Wongphechauxsorn et al., 2024). Zhao et al. (2020) identified 110 compact radio sources, and Zhao et al. (2022) discovered 64 hypercompact sources in the same area, some of which may be radio pulsars. Multi-wavelength analyses have provided indirect constraints on the pulsar population. For instance, Wharton et al. (2012) estimated that the inner parsec (i.e. a region $\approx$300 times smaller than the $\gamma$-ray excess area) around the Galactic centre could host up to $10^{3}$ active pulsars. Over the years, several studies have attempted to estimate the size of the MSP population that could be responsible for the $\gamma$-ray excess (Hooper and Linden, 2016; Ploeg et al., 2017; Dinsmore and Slatyer, 2022). A recent study by Holst and Hooper (2025) derived a $\gamma$-ray luminosity function using field MSPs and concluded that if $\gamma$-ray luminosity remains constant over time, the Galactic bulge could host as many as $3.6\times 10^{4}$ MSPs. If instead the MSP $\gamma$-ray luminosity decays with time, the estimated population size rises to $2\times 10^{5}$ MSPs.
@@
The NSs and MSPs are also promising sources of continuous gravitational-wave emission. If a rapidly spinning NS has even a small amount of ellipticity—either imprinted at birth, caused by starquakes, accretion from a companion, or sustained by internal magnetic fields—it will be characterised by a non-zero mass quadrupole. This leads to the emission of gravitational waves at a frequency twice the star’s rotational frequency, typically in the tens to thousands of hertz range, which falls within the sensitivity band of ground-based detectors (Friedman and Schutz, 1978). However, these continuous signals are intrinsically weak, and at the distance of the Galactic centre, they remain well below the detection threshold of current detectors (e.g. Abbott et al., 2022; Riles, 2023; Steltner et al., 2023). Next-generation detectors, such as the Einstein Telescope and Cosmic Explorer, are expected to achieve the sensitivity required to probe these sources in the Galactic centre (e.g. Pagliaro et al., 2023; Bartel and Profumo, 2024).
@@
For NSs and MSPs in binaries with orbital periods less than about two hours, the gravitational-wave signal lies in the millihertz band and is accessible to space-based detectors, such as the Laser Interferometer Space Antenna (LISA; Colpi et al., 2024), even at distances comparable to the Galactic centre and beyond (Seto, 2019; Lau et al., 2020; Korol and Safarzadeh, 2021; Wagg et al., 2022). In this article, we investigate whether LISA can probe the MSP population in binary systems in the Galactic bulge and thus provide an independent test of their proposed contribution to the $\gamma$-ray excess. We argue that because gravitational-wave detection is subject to entirely different selection effects than electromagnetic searches for MSPs, LISA offers a promising new way to test the MSP origin of the $\gamma$-ray excess. To illustrate this potential, we present the first forecasts for MSP binaries in the Galactic bulge detectable by LISA.
@@
This article is structured as follows. In Section 2, we briefly review the literature on the formation of MSP binaries and summarise their gravitational-wave detectability with LISA. In Section 3, we describe the assumptions and methodology behind our population synthesis of the bulge. In Section 4, we present our forecasts for different bulge formation scenarios and discuss the likelihood of detecting and characterising the MSP population with LISA. Finally, in Section 5, we discuss our findings in a broader context, and we summarise our conclusions in Section 6.
@@ 2 Galactic millisecond pulsar binary population
The Australia Telescope National Facility (ATNF) pulsar catalogue v 2.6.1111https://www.atnf.csiro.au/research/pulsar/psrcat (Manchester et al., 2005) contains information on 624 pulsars with spin periods shorter than $P\leq 30$ ms, most of which are classified as MSPs. A significant fraction of these pulsars are located in globular clusters. For example, the most recent compilation by Paulo Freire [^2] https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html summarises properties of 345 pulsars in 45 clusters, indicating that at least 55% of all known MSPs reside in globular clusters. Whether formed dynamically in globular clusters or through isolated binary evolution in the field, MSPs can end up in a variety of binary configurations. Below, we introduce those configurations that are mentioned in this study, drawing primarily from the classification and discussion presented in Tauris and van den Heuvel (2023).
@@
Many MSPs are thought to descend from LMXBs, in which a NS accretes material from a companion star and is spun up to millisecond spin periods. A large fraction of known MSPs are found in binaries with helium-core white dwarf (He WD) companions. Other systems host more massive carbon–oxygen-core WDs (CO WDs) or oxygen–neon–magnesium-core WDs (ONeMg WDs). Some MSPs reside in tight ultra-compact X-ray binaries (UCXBs) with orbital periods as low as a several minutes, while others are found in eclipsing systems known as spiders. These are typically categorised as black widows, with extremely low-mass companions (<0.1<0.1 M), or redbacks, with more massive companions (0.10.10.40.4 M). Most recently, Yang et al. (2025) discovered an MSP with a massive companion (up to $1.6$ M), most likely a stripped helium star (He star).
@@
Figure 1: Orbital periods and companion masses for MSPs compiled from the ATNF Pulsar Catalogue v2.6.0. The companion mass shown is the median mass, computed assuming an orbital inclination of $i=60^{\circ}$. All data points (blue cross symbols) correspond to radio-detected MSPs. Among them, $\gamma$-ray sources are highlighted with green downward-pointing triangles, and those located in globular clusters (GCs) are marked with orange upward-pointing triangles. Grey error bars indicate lower and upper limits on WD companion masses. For reference, the theoretical orbital period–companion mass relation from Tauris and Savonije (1999) for Population I stars is shown as a solid red line. The minimum gravitational-wave frequency detectable by LISA is shown as a dashed black line.
@@
Figure 2: Top panel: Distribution of companion masses for the MSP sample. The solid red lines show a power-law distribution with a slope of 4-4, plotted for comparison. Bottom panel: Distribution of binary orbital periods for the MSP sample. Overplotted are the estimated pulsar detection efficiencies in radio surveys, based on the modelling by Bagchi et al. (2013). Two search strategies are shown: $\gamma_{1}$ for a standard (non-accelerated) search, and $\gamma_{2}$ for an acceleration search. These detection efficiency curves are computed assuming the following orbital parameters: eccentricity $e=0$, $M_{\mathrm{MSP}}=1.4$ M, integration time $T\mathrm{obs}=1000$ s, WD companion mass derived from the Tauris and Savonije (1999) relation, spin period of 3 ms, orbital inclination $i=60^{\circ}$, and longitude of periastron $\Omega=30^{\circ}$.
@@ 2.1 MSPs formation and orbital parameters
Tauris and Savonije (1999) identified a relation between the mass of a He WD and the orbital period at the end of stable mass transfer. This relation arises because stable mass transfer during the LMXB stage uniquely determines the orbital separation based on the donor star’s density. The relation is given by
@@
To better understand the observed MSP population, we selected all millisecond radio pulsars from the ATNF pulsar catalogue. We selected sources using the commonly adopted definition with $P_{\rm s}<30$ ms and $\dot{P}_{\rm s}<10^{-19}$ s/s. We also retrieved information about associations with globular clusters and the presence of the $\gamma$-ray emission. The resulting distribution of orbital periods and companion masses is shown in Fig. 1. The MSPs fall broadly into two groups: (1) binaries with He WD companions that follow the Tauris and Savonije (1999) relation (solid red line) and (2) binaries with low-mass companions—including systems that are the products of UCXB evolution, redbacks, and black widows. The ATNF catalogue also includes some MSPs likely formed through a common-envelope channel with CO WD companions, as well as rare systems in which MSPs are paired with NSs (i.e. those with companion masses >0.5>0.5 M in Fig. 1). Importantly, these latter systems are only partially recycled, meaning they have longer spin periods and higher magnetic fields. As a result, their spin-down luminosities are typically too low to produce detectable $\gamma$-ray emission, and these pulsars are therefore not seen as $\gamma$-ray sources. Nevertheless, some of them may be prominent gravitational-wave sources because of their larger component masses.
In Fig. 2 we show the distributions of companion masses (top) and orbital periods (bottom) for the MSPs in our ATNF catalogue selection. The companion mass distribution can be approximated by a power law with an exponent of 4-4, which matches the observed trend down to $\sim 0.2$ M, particularly for systems in globular clusters (orange distribution). In the right panel, we compare the orbital period distribution with the detection efficiency curves from Bagchi et al. (2013) for standard ($\gamma_{1}$) and acceleration ($\gamma_{2}$) radio searches. The number of known MSPs drops sharply at short orbital periods, closely following the shape of the $\gamma_{1}$ efficiency curve. This suggests that observational biases significantly affect the detection of binaries with orbital periods shorter than one day, and implies that the intrinsic population may be more uniformly distributed in logarithmic orbital period space than the observed sample indicates.